# Search for Axion dark matter with the QUAX–LNF tunable haloscope

### v1

@@ Acknowledgments
This work is supported by INFN (QUAX experiment), by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under Contract No. DE-AC02-07CH11359, by the European Union’s FET Open SUPERGALAX project, Grant No. 863313, and by the PNRR MUR project number PE0000023-NQSTI.
@@
The authors wish to acknowledge the technical support of Anna Calanca, Marco Beatrici, Stefano Lauciani, Maurizio Gatta, Giuseppe Pileggi, Giuseppe Papalino and Daniele Di Bari, as well as of the Mechanical Workshop Department of the Mechanical Design and Construction Service (SPCM) of LNF, and of the National Laboratories of Legnaro for electropolishing the inner surface of the OFHC cavity.
@@ V.1 Dark Photons
$\chi=g_{a\gamma\gamma}\frac{B}{m_{\scriptscriptstyle\text{DP}}\left|\cos\theta\right|}$ (9)

### v3

@@ Acknowledgments
This work is supported by INFN (QUAX experiment), by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under Contract No. DE-AC02-07CH11359, by the European Union’s FET Open SUPERGALAX project, Grant No. 863313, and by the PNRR MUR project number PE0000023-NQSTI. This research was also supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
@@
The authors wish to acknowledge the technical support of Anna Calanca, Marco Beatrici, Stefano Lauciani, Maurizio Gatta, Giuseppe Pileggi, Giuseppe Papalino and Daniele Di Bari, as well as of the Mechanical Workshop Department of the Mechanical Design and Construction Service (SPCM) of LNF, and Fabrizio Stivanello and Eduard Chyhyrynets from the National Laboratories of Legnaro for electropolishing the inner surface of the OFHC cavity.
@@ V.1 Dark Photons
$\chi=g_{a\gamma\gamma}\frac{B}{m_{\scriptscriptstyle\text{DP}}\left|\cos\theta\right|}\,,$ (9)
@@
V.2 Calibration procedure
To calibrate the output power, we need to know the gain of the readout line ($L5$). We follow the procedure developed in [66]. The rf diagram of Fig. 1 can be represented, in dB units, by the following system of three coupled equations:
$S_{51}=G_{L1}+S_{21}^{\scriptscriptstyle\text{cav}}(\nu_{c},Q_{0},\beta)+G_{L5},$
$S_{53}=G_{L3}+S_{22}^{\scriptscriptstyle\text{cav}}(\nu_{c},Q_{0},\beta)+G_{L5},$ (10)
$S_{13}=G_{L1}+S_{12}^{\scriptscriptstyle\text{cav}}(\nu_{c},Q_{0},\beta)+G_{L3},$
where $S_{51}$, $S_{53}$ and $S_{13}$ are the measured scattering parameters, $G_{L1}$, $G_{L3}$ and $G_{L5}$ are the attenuations/gains of the respective input and output lines, and $S_{ij}^{\scriptscriptstyle\text{cav}}$ are the Lorentzian functions of the cavity (here given in linear form):
$S_{21}^{\scriptscriptstyle\text{cav}}=S_{12}^{\scriptscriptstyle\text{cav}}=\left\lvert\frac{2\sqrt{\beta_{1}\beta_{2}}}{1+\beta_{1}+\beta_{2}+jQ_{0}\delta}\right\rvert,$ (11)
$\begin{split}S_{22}^{\scriptscriptstyle\text{cav}}=\,\,&\bigg{\lvert}\frac{\beta_{2}^{2}-(1+\beta_{1})^{2}-(Q_{0}\delta+q)^{2}}{(1+\beta_{1}+\beta_{2})^{2}+Q_{0}^{2}\delta^{2}}+\\ &-j\frac{2\beta_{2}Q_{0}\delta}{(1+\beta_{1}+\beta_{2})^{2}+Q_{0}^{2}\delta^{2}}\bigg{\rvert},\end{split}$ (12)
where $\beta_{2}$ is the coupling to the mobile antenna, $\delta$ is equal to $\omega/\omega_{0}-\omega_{0}/\omega$, and $q$ is a parameter accounting for the small asymmetry in the reflection coefficient.
At each frequency step in Table 1 we perform a simultaneous fit of the $S_{51}$, $S_{53}$ and $S_{13}$ data. Note that in the fits $\beta_{1}$ is kept fixed at $1.4\times 10^{-3}$, the value expected from simulations, since it was verified to be less than $7\times 10^{-3}$ combining measurements of $L1$ and $L2$ data.
Upon solving Eqs. (10) using the parameters found from the fit, the gain of the readout line is
$G_{L5}=\frac{(S_{51}-S_{21}^{\scriptscriptstyle\text{cav}})+(S_{53}-S_{22}^{\scriptscriptstyle\text{cav}})-(S_{13}-S_{12}^{\scriptscriptstyle\text{cav}})}{2},$ (13)
evaluated as the average of the spectrum. The maximum spread of the $G_{L5}$ values within all the subruns is only 0.4 dB, so that we can reasonably consider the gain constant in the evaluated frequency range and equal to the average value of $G_{L5}=70.62$ dB.
This value is referred at the input of the splitter. Since the $S$ parameters are measured to and from the VNA ports, to calculate the right value for $G_{L5}$ the contributions of the splitter ($-3$ dB) and of the cable from the splitter to the VNA have been subtracted. Then, a calibration of the downconversion electronics from the splitter input to the DAQ is also necessary. This has been performed only once and is valid for all subruns, since it does not depend on the frequency change. We send a known power at the splitter and measure the downconverted signal at the DAQ input with a Spectrum Analyzer, which gives an absolute power. We repeat the measurement at different LO frequencies to cover all the acquisition bandwidth of the ADC board, from $-1$ to 1 MHz. The calibration curve is shown in Fig. 9, where the roll-off is due to the ADC internal filters.
Finally, the downconversion calibration curve and the gain $G_{L5}$ are subtracted to the raw power spectra acquired by the ADC to obtain calibrated power spectra (as the one in Fig. 5), where the power level is therefore referred at the cavity readout port.
Figure 9: Calibration of the downconversion electronics in the frequency band of the ADC.